Marrisa Anderson is a huge fan of the science show "Mr. Wizard" so she decides to make her own knock-off skit of it called "Ms. Wizard". Because of demand it is held at a large auditorium. Tickets are sold at the following prices, Children are 4.50$ Adults are 9.75$ and Seniors are 5.50$.There is a raffle she is doing where she gives away the whole show on a DVD. Each child receives one raffle ticket, seniors receive 2 and adults receive 3.
The show is a huge success and only on audience helpers catches on fire. After the show Marrisa checks the number of people that attended and finds out that 1,118 people attended the science spectacular. She counts the total amount of money she made and finds out she made $8,037.50. The number of raffle tickets total is 2,518. She wants to find out the following, how many science loving people attending that were children, how many were adults and how many were seniors?
(Note the strategy known as "Guess and Check" is not allowed to solve this problem)

Answer by CubeyThePenguin(3113)

a = adult's tickets
c = children's tickets
s = senior's tickets

a + c + s = 1118
9.75a + 4.5c + 5.5s = 8037.5
3a + c + 2s = 2518

You can subtract equation 3 from equation 1 to eliminate the variable "c" and then proceed from there.